# CosForce: A Force-Based General Pedestrian Model with Anticipation and Reaction Mechanisms

### v3

@@ Abstract
In this study, we developed a force-based general pedestrian model named CosForce. To the best of our knowledge, this may represent the simplest version of the force-based method. The model employs cosine functions to characterize asymmetric interactions, implicitly incorporating anticipation and reaction mechanisms. By focusing on binary interactions, the CosForce model provides new insights into pedestrian modeling while achieving linear computational complexity. Two specific scenarios in crowd dynamics were analyzed: self-organization (entropy decrease) and crowd collapse (entropy increase). The average normalized speed and order parameter were introduced to quantitatively describe the processes of crowd dynamics. Quantitative evaluations demonstrate that phase separation in crowds is effectively reproduced by the model, including lane formation, stripe formation, and cross-channel formation. Next, in the simulation of mass gathering, within a density-accumulating scenario, processes of critical phase transition in high-density crowds are clearly revealed through time series observations of the order parameter. These findings provide valuable insights into crowd dynamics.
Force-based models, Crowd dynamics, Phase separation, Critical transitions, Numerical simulation
@@ 1 Introduction
The study of crowd dynamics has evolved over several decades. Although pedestrian dynamics are closely related to our daily lives, our understanding of crowd behavior remains incomplete. As a field at the intersection of traffic dynamics and collective dynamics, it has naturally integrated perspectives and theories from various disciplines, such as fundamental diagram theory (Zhang et al., 2012) and critical transition mechanisms (Scheffer et al., 2012; Vicsek et al., 1995; Szabo et al., 2006), among others. Empirical studies of Hajj (Helbing et al., 2007), the Love Parade disaster (Ma et al., 2013), and the San Fermín festival (Gu et al., 2024) have demonstrated the fluid-like properties of dense crowds. These findings indicate that zero-flow density does not strictly exist in crowd systems, establishing a new paradigm for crowd analysis. However, a comprehensive theoretical framework that fully captures crowd dynamics is still lacking. Empirical studies provide extensive data on pedestrian behavior under normal conditions, yet precise data for extreme scenarios remain scarce (Haghani & Sarvi, 2018). Moreover, due to the self-driven and social nature of crowds, behavioral diversity makes it challenging to establish universally applicable observations. These difficulties complicate the simplification of crowd movement into a straightforward particle-interaction process for analysis, yet such an approach remains essential.
@@
Microscopic pedestrian dynamics models mainly focus on real-time interactions based on operational-level simulation, such as the cellular automaton model (CA) (Burstedde et al., 2001), the force-based model (Helbing & Molnar, 1995; Chraibi et al., 2010), the reciprocal velocity obstacles model (Van den Berg et al., 2008), and the heuristic model (Moussaïd et al., 2011; Xiao et al., 2016), have been widely applied. Researchers have also sought inspiration from interdisciplinary knowledge, such as the concept of order parameters derived from Landau theory, the mean-field game theory (Bonnemain et al., 2023), and the PLE model (Guy et al., 2010) based on Principle of Least Action (Alahi et al., 2016; Gupta et al., 2018). Based on the memoryless property of Markov processes, physics-based approaches incorporate stochastic noise to avoid deterministic simulation. Another category, trajectory prediction methods based on LSTM architecture, from a statistical perspective, where time-series inference is elevated to a paramount position. Numerical validation results suggest that data-driven models may surpass physics-based approaches, at least in the tasks of short-term pedestrian trajectory prediction (a few seconds, low-density). Nevertheless, in analyzing crowd dynamics, physics-based modeling remains the most reliable approach as it operationalizes the core principles of human reasoning.
@@
Under normal conditions, pedestrians exhibit noncontact interactions, where collision avoidance behavior is constrained by empirical relationships. The most commonly used principle is the linear relationship between headway (which can be approximated in a 2D space by the nearest neighbor in the direction of motion) and speed. Researchers can add new rules to investigate specific mechanisms (Liang et al., 2021; Shang et al., 2024). Among these, anticipation and reaction have been extensively analyzed. Collision avoidance rules can be introduced as modeling based on anticipation behavior (Zanlungo et al., 2011; Gerlee et al., 2017; Lü et al., 2020; Hu & Chen, 2024; Xu et al., 2021). Under extreme density, the collision avoidance mechanisms of pedestrians completely fail, and Newton’s third law becomes the fundamental principle governing crowd dynamics. The motion of the crowd therefore displays properties similar to particle hydrodynamics (Van Toll et al., 2020). Force-based models provide a comprehensive physical framework for granular simulation, with the capability to accurately model interactions within dense crowds. Among these, a potential issue is that the complex models often perform poorly in cross-scenario simulations. Such models are frequently effective in specific scenarios; however, the optimization for particular objectives may, in fact, undermine their generalization capability. More importantly, the introduction of additional components complicates the relatively simple principle of pedestrian dynamics. The cumbersome rules and formulas limit the ability of model for provide straightforward insight in pedestrian dynamics. Furthermore, Simulation results are challenging to validate objectively. In this regard, it may be more effective to develop a general model by simplifying models rather than by adding complexity. Everyone naturally engages with motion in their daily lives. Therefore, preserving empirical insights and intuitive understanding is essential. This principle, as discussed below, forms the cornerstone of the model proposed in this paper.
@@
Empirical observations of pedestrian behavior reveal that individuals primarily focus on the dynamics in front of them, largely disregarding those behind. Prior to introducing the model, the Horizontal Field of Attention (HFA, $\mathbf{\Omega}\in{\mathbb{R}^{2}}$) in our model is defined. The HFA is constrained by the attentional eccentricity angle $\phi$ and the attentional depth $h$, as depicted in Fig.1. The angle $\phi$ represents the angle between the boundary of the sector-shaped HFA and the pedestrian’s direction of movement. The attentional depth $h$ can theoretically extend to infinity. To minimize computation, $h$ can be set to the minimum headway allowing pedestrians to maintain free motion.
@@
The models based on collision avoidance algorithms has proven to be highly effective in studying interactions within crowds and is currently widely used in modeling two-dimensional multi-agent systems for obstacle avoidance (Van den Berg et al., 2008; Tordeux et al., 2016). A potential issue is that the precise collision avoidance algorithm actually limits its capability for high-accuracy crowd simulation: in crowds, collisions are common and, in some situations, even inevitable. This discrepancy arises mainly because pedestrian speed and orientation are updated based on empirical estimates rather than precise numerical calculations. In reality, pedestrians can not perform such calculations, their motion mainly rely on approximations. In this consideration, the collision process is crucial for pedestrian simulation, especially in dense crowds.
@@
Research based on collective dynamics typically establishes models from a holistic perspective, wherein macro-level rules govern the evolution of groups, such as the BOID model (Reynolds, 1987) and network models (Bode et al., 2011; Allen et al., 2017). Based on metrics (Vicsek et al., 1995), topological (Ballerini et al., 2008), or visual connection (Rosenthal et al., 2015; Wirth et al., 2023), the state of individuals exhibits a strong correlation with the characteristics of their surrounding neighborhoods (coherence of motion). The complexity and variability of pedestrian dynamics hinder the approximation of crowds to fish schools or flocks of birds. However, the paradigms of interactions observed can indeed be transferred across species.
@@
In highly stochastic and dynamic crowds, pedestrians can realistically focus on only a limited number of targets, typically their nearest neighbors within their attention field. From this perspective, models that employ binary interactions (arc connection with a betweenness of 1) rather than metric-based interactions may be more concise and efficient.
@@
The phenomenon of self-organization among pedestrians is characterized by entropy decrease and the emergence of steady states within crowd systems. From a reductionist perspective, we conjecture that such entropy decrease is presumably induced by local cooperative interactions among conflicting pedestrians (Rand et al., 2014; Su et al., 2022; Bonnemain et al., 2023; Zablotsky et al., 2024). In a typical crossing flow scenario, as depicted in Fig.2(a), there is a potential risk of collision between crossing pedestrians. Anticipation mechanism, involving a negotiation or game process, can facilitate "phase separation" between pedestrians moving in different directions. The interaction results between pedestrians from distinct directions can be represented by a simple payoff matrix, as shown in Fig.2(b). Cooperative games contribute to entropy decrease within the system, leading to the emergence of spontaneous order, as illustrated in Fig.2(c). In this consideration, the local cooperation rule is implicitly embedded in our model through asymmetric force interactions governed by the cosine constraint.
From the perspective of applicability and simplicity, the general pedestrian model of the operational layer should incorporate the following characteristics: Empirical foundations: grounded in empirical data and aligned with our experience; Minimalist principles: employing the simplest rules to reveal core mechanisms while allowing straightforward analysis; Diversity in model performance: parameters can transfer across scenarios, consistent with fundamental diagrams, diversity in simulation (e.g., phenomena of pedestrian self-organization); Computational efficiency: lower computational cost compared to other models in the same category. To this end, based on the above requirements, a general model of pedestrian dynamics is established in this paper. The rules of the model, developed from first principles, were further refined in accordance with the classic social force model to enhance simulation accuracy. Additionally, this minimalist approach facilitates the analytical examination of pedestrian interactions. The subsequent sections of the paper unfolded as follows: In section 2, considering the anticipation and reaction behavior of pedestrians, we established the CosForce model. In section 3, empirical validations were conducted for single-file motion and unidirectional flow. Subsequently, to investigate two specific mechanisms in crowd dynamics, namely the spontaneous processes of entropy decrease and entropy increase, we introduced fundamental parameters for quantitative evaluation the self-organization processes in section 4 and the phenomenon of crowd collapse in section 5. Based on the modeling assumption of binary interactions, we validated the advantage of linear time complexity in computational efficiency in section 6 and discussed the limitations. Finally, section 7 summarized the conclusions drawn from the study.
@@ 2.1 Constraint of Space-speed Relationship
The correlation between speed and space has been thoroughly investigated, as depicted in Fig.2. Under the assumption of isotropic pedestrian interactions, it has been demonstrated that, in constrained scenarios, velocity is inversely correlated with local density (i.e., the speed–density relationship), as articulated in Eq.1, with a constant parameter denoted as $k$ (dimension specified as $\mathrm{m}^{-1}\,\mathrm{s}^{-1}$). This empirical relationship serves as the foundation within force-based models, wherein the variation in pedestrian force concerning spatial metrics is characterized by a power function, as elucidated in Eq.2, $\tau$ represents the relaxation time.
@@
[…]
Furthermore, in one-dimensional motion, a linear empirical relationship between the headway and speed is typically employed, often associated with the principle of constant time headway. Considering the anisotropy of the pedestrian movement space, this linear relationship, based on the nearest neighbor distance, can similarly be extended to two-dimensional motion, as illustrated in Eq.3. In this context, a linear function can be employed to describe the variation of pedestrian force with respect to the relative distance of nearest neighbor, as shown in Eq.4.
@@
[…]
@@
2.2 Modeling of Anticipation and Reaction Behaviors
Symmetric and Aymmetric forces
Figure 4: Schematic of nearest neighbor interaction: (a) reciprocal interaction (anticipation mechanism), (b) non-reciprocal interaction (reaction mechanism).
@@ 2.2 Modeling of Anticipation and Reaction Behaviors
For the convenience of differential computation, the tanh function can be employed to describe its continuous form, which aligns with the formulation used to represent the desired velocity in traffic flow. The continuous form is provided in Eq.9. Fig.5 illustrated the influence of the nearest-neighbor relative distance $d$ and angle $\theta$ on the force term.
@@ 2.3 Model Rules
Based on the framework of a force-based model, the model is characterized by a self-driven force $\mathbf{f}_{i}$, a repulsive force $\mathbf{f}_{ij}$ exerted by the nearest entity (pedestrian or wall $j$, $j\in\mathbf{\Omega}:=\left\|{{\mathbf{d}_{ij}}}\right\|<h\wedge\angle\left({{\mathbf{v}_{i}},{\mathbf{d}_{ij}}}\right)<\phi$), and the contact forces $\mathbf{f}_{c}$ that arise during collisions. The global constraints can be expressed as follows:
$\mathbf{f}=\mathbf{f}_{i}+\mathbf{f}_{ij}+\sum_{\|\mathbf{d}_{ij}\|<r_{ij}}\mathbf{f}_{c}=m\mathbf{a}$ (10)
@@
Here, $\theta$ is the angle between the relative distance and the relative velocity, $\theta=\angle\left(\mathbf{v}_{ij},\mathbf{d}_{ij}\right)\in\left[0,\pi\right]$. Other symbols are defined as follows: $r_{ij}=r_{i}+r_{j}$, $\mathbf{v}_{ij}=\mathbf{v}_{i}-\mathbf{v}_{j}$, and $\mathbf{d}_{ij}=\mathbf{x}_{j}-\mathbf{x}_{i}$, ${\mathbf{n}_{ij}}=-{\mathbf{d}_{ij}}/\left\|{{\mathbf{d}_{ij}}}\right\|$. In this process, obstacles are modeled as stationary pedestrians and are therefore included in the calculation of the repulsive force without requiring any additional rules. Different from the point-to-point interaction between pedestrians, the interaction between pedestrians and obstacles follows a point-to-line relationship. Considering all potential collision scenarios, the attentional eccentricity angle $\phi$ for pedestrian-wall interaction is set to a constant value of $\pi/2$.
@@
As mentioned earlier, the pedestrian is represented as a compressible particle, described by an exponential function. When the pedestrian’s body undergoes compression, the resulting repulsive force increases exponentially with the compression magnitude. A fully repulsive core exists as the compression limit for the pedestrian particle. The parameter settings are presented in Tab.1, where $m=20$ kg, $r=0.2$ m, $\tau=0.5$ s, $t_{h}=1.3$ s, $\lambda=0.8$ and $v_{\rm{max}}$= 1.4 m/s are global parameters that remain constant across all simulations. The variable parameters $\phi$, and $\alpha$ will be adjusted as required. The frame rate for all simulations is set as 30, corresponding to a difference time of $1/30$ s.
@@
Parameter Description Reference value
$m$ Mass 60 kg
$v_{\rm{max}}$ Maximum speed 1.4 m/s
$r$ Radius 0.2 m
$\tau$ Relaxation time 0.5 s
$t_{h}$ Approximate time headway 1.3 s
$\phi$ Attentional eccentricity angle [0 - $\pi$] rad
$\lambda$ Dimensionless coefficient 0.8/0.02 (scale in centimeters/meters)
$\alpha$ Dimensionless coefficient [0-1]

### v4

@@ Abstract
This paper develops a general force-based pedestrian model named CosForce, in which cosine functions are employed to describe asymmetric interactions. These functions implicitly capture the mechanisms of anticipation and reaction. By focusing on binary interactions, the CosForce model provides new insights into pedestrian modeling while achieving property of linear time complexity. Two specific scenarios in crowd dynamics were analyzed: self-organization (entropy decrease) and crowd collapse (entropy increase). The average normalized speed and order parameter were introduced to quantitatively describe the processes of crowd dynamics. Quantitative evaluations demonstrate that phase separation in crowds is effectively reproduced by the model, including lane formation, stripe formation, and cross-channel formation. Next, in the simulation of mass gathering, within a density-accumulating scenario, processes of critical transition in high-density crowds are clearly revealed through time series observations of the order parameter. These findings provide valuable insights into crowd dynamics.
Force-based method, Crowd dynamics, Phase separation, Critical transitions, Numerical simulation
@@ 1 Introduction
The study of crowd dynamics has evolved over several decades. Although pedestrian dynamics are closely related to daily lives, our understanding of crowd behavior remains incomplete. As a field at the intersection of traffic dynamics and collective dynamics, it has naturally integrated perspectives and theories from various disciplines, such as theory of pedestrian fundamental diagram (Seyfried et al., 2005; Zhang et al., 2012) and critical transition mechanisms (Scheffer et al., 2012; Vicsek et al., 1995; Szabo et al., 2006), among others. Field studies of Hajj (Helbing et al., 2007), the Love Parade disaster (Ma et al., 2013), and the San Fermín festival (Gu et al., 2024) have demonstrated the fluid-like properties of dense crowds. These findings indicate that zero-flow density does not strictly exist in crowd systems, establishing a new paradigm for crowd analysis. However, a comprehensive theoretical framework that fully captures crowd dynamics is still lacking. Empirical studies provide extensive data on pedestrian behavior under normal conditions, yet precise data for extreme scenarios remain scarce (Haghani & Sarvi, 2018). Moreover, due to the property of self-propelled and social nature of crowds, behavioral diversity makes it challenging to establish universally applicable observations. These difficulties complicate the simplification of crowd movement into a straightforward particle-interaction process for analysis, yet such an approach remains essential.
@@
Microscopic pedestrian dynamics models mainly focus on real-time interactions based on operational-level simulation, such as the cellular automaton model (CA) (Burstedde et al., 2001), the force-based model (Helbing & Molnar, 1995; Chraibi et al., 2010), the collision avoidance model (Van den Berg et al., 2008; Tordeux et al., 2016), and the heuristic model (Moussa¨ıd et al., 2011), have been widely applied. Researchers have also sought inspiration from interdisciplinary knowledge, such as the concept of order parameters derived from Landau theory, the mean-field game theory (Bonnemain et al., 2023), and the PLE model (Guy et al., 2010) based on Principle of Least Action. Based on the memoryless property of Markov processes, physics-based approaches incorporate stochastic noise to avoid deterministic simulation. Another category, trajectory prediction methods, from a statistical perspective, where time-series inference is elevated to a paramount position (Alahi et al., 2016). Numerical validation results suggest that data-driven models may surpass physics-based approaches, at least in the tasks of short-term pedestrian trajectory prediction (a few seconds, low-density). Nevertheless, in analyzing crowd dynamics, physics-based modeling remains the most reliable approach as it operationalizes the core principles of human reasoning.
@@
Under normal conditions, pedestrians exhibit noncontact interactions, where collision avoidance behavior is constrained by empirical relationships. The most commonly used principle is the linear relationship between headway (which can be approximated in a 2D space by the distance with nearest neighbor in the direction of motion) and speed. Researchers can add new rules to investigate specific mechanisms (Liang et al., 2021; Shang et al., 2024). Among these, anticipation and reaction have been extensively analyzed. Collision avoidance rules can be introduced as modeling based on anticipation behavior (Zanlungo et al., 2011; Gerlee et al., 2017; Lü et al., 2020; Hu & Chen, 2024; Xu et al., 2021). In extreme density, the collision avoidance mechanisms of pedestrians will completely fail, and Newton’s third law becomes the fundamental principle governing crowd dynamics. The motion of the crowd therefore displays properties similar to particle hydrodynamics (Van Toll et al., 2020). Force-based models provide a comprehensive physical framework for particle simulation, with the capability to accurately model interactions within dense crowds. Among these, a potential issue is that the complex models often perform poorly in cross-scenario simulations. Such models are frequently effective in specific tasks; however, the optimization for particular objectives may, in fact, undermine their generalization capability. More importantly, the cumbersome rules and formulas limit the ability of model for provide straightforward insight in pedestrian dynamics. Furthermore, Simulation results are challenging to validate objectively. In this regard, it may be more effective to develop a general model by simplifying models rather than by adding complexity. Everyone naturally engages with motion in their daily lives. Therefore, preserving empirical insights and intuitive understanding is essential. This principle, as discussed below, forms the cornerstone of the model proposed in this paper.
@@
Empirical observations of pedestrian behavior reveal that individuals primarily focus on the dynamics in front of them, largely disregarding those behind. Prior to introducing the model, the Horizontal Field of Attention (HFA, $\mathbf{\Omega}\in{\mathbb{R}^{2}}$) in our model is defined. The HFA is constrained by the attentional eccentricity angle $\phi$ and the attentional depth $h$, as depicted in Fig.1. The angle $\phi$ represents the angle between the boundary of the sector-shaped HFA and the pedestrian’s direction. The attentional depth $h$ can theoretically extend to infinity. To minimize computation, $h$ can be set to the minimum headway allowing pedestrians to maintain free motion.
@@
The models based on collision avoidance algorithms has proven to be highly effective in studying interactions within crowds and is currently widely used in modeling 2-D multi-agent systems for obstacle avoidance (Van den Berg et al., 2008; Tordeux et al., 2016). A potential issue is that the precise collision avoidance algorithm actually limits its capability for high-accuracy crowd simulation: in crowds, collisions are common and, in some situations, even inevitable. This discrepancy arises mainly because pedestrian speed and orientation are updated based on empirical estimates rather than precise numerical calculations. In reality, pedestrians can not perform such calculations, their motion mainly rely on approximations. In this consideration, the collision process is crucial for pedestrian simulation, especially in dense crowds.
@@
Research based on collective dynamics typically establishes models from a bottom-up perspective, wherein micro-level rules govern the evolution of groups, such as the BOID model (Reynolds, 1987) and network models (Bode et al., 2011; Allen et al., 2017). Based on metrics (Vicsek et al., 1995), topological (Ballerini et al., 2008), or visual connection (Rosenthal et al., 2015; Wirth et al., 2023), the state of individuals exhibits a strong correlation with the characteristics of their surrounding neighborhoods (coherence of motion). The complexity and variability of pedestrian dynamics hinder the approximation of crowds to fish schools or flocks of birds. However, the paradigms of interactions observed can indeed be transferred across species.
@@
In highly stochastic and dynamic crowds, pedestrians can realistically focus on only a limited number of targets, typically their nearest neighbors within their attention field. From this perspective, models that employ binary interactions rather than metric-based or pairwise interactions may be more concise and efficient.
@@
The phenomenon of self-organization among pedestrians is characterized by entropy decrease and the emergence of steady states within crowd systems. From a reductionist perspective, we conjecture that such entropy decrease is presumably induced by local cooperative interactions among conflicting pedestrians (Rand et al., 2014; Su et al., 2022; Bonnemain et al., 2023; Zablotsky et al., 2024). In a typical crossing flow scenario, as depicted in Fig.2(a), there is a potential risk of collision between crossing pedestrians. Anticipation mechanism, involving a negotiation or game process, can facilitate "phase separation" between pedestrians moving in different directions. The interaction results between pedestrians from distinct directions can be represented by a simple payoff matrix, as shown in Fig.2(b). Cooperative games contribute to entropy decrease within the system, leading to the emergence of spontaneous order, as illustrated in Fig.2(d). In this consideration, the local cooperation rule is implicitly embedded in our model through asymmetric force interactions governed by the cosine function.
From the perspective of applicability and simplicity, the general pedestrian model of the operational layer should incorporate the following characteristics: Empirical foundations: grounded in empirical data and aligned with our experience; Minimalist principles: employing the simplest rules to reveal core mechanisms while allowing straightforward analysis; Diversity in model performance: parameters can transfer across scenarios, consistent with fundamental diagrams, diversity in simulation (e.g., phenomena of pedestrian self-organization); Computational efficiency: lower computational cost compared to other models in the same category. To this end, based on the above requirements, a general model of pedestrian dynamics is established in this paper. The rules of the model, developed from minimalism principles, were further refined in accordance with the classic social force model to enhance simulation accuracy. Additionally, this minimalist approach facilitates the analytical examination of pedestrian interactions. The subsequent sections of the paper unfolded as follows: In section 2, considering the anticipation and reaction behavior of pedestrians, we established the CosForce model. In section 3, empirical validations were conducted for single-file motion and unidirectional flow. Subsequently, to investigate two specific mechanisms in crowd dynamics, namely the spontaneous processes of entropy decrease and entropy increase, we introduced fundamental parameters for quantitative evaluation the self-organization processes in section 4 and the phenomenon of crowd collapse in section 5. Based on the modeling assumption of binary interactions, we validated the advantage of linear time complexity in computational efficiency in section 6 and discussed the limitations. Finally, section 7 summarized the conclusions drawn from the study.
@@ 2.1 Constraint of Space-speed Relationship
The correlation between speed and space has been thoroughly investigated in experimental studies, as depicted in Fig.3. Under the assumption of isotropic pedestrian interactions, it has been demonstrated that, in constrained scenarios, velocity is inversely correlated with local density (i.e., the speed–density relationship), as articulated in Eq.1, with a constant parameter denoted as $k$ (dimension specified as $\mathrm{m}^{-1}\,\mathrm{s}^{-1}$). This empirical relationship forms the foundation of force-based methods, treating the relaxation time as a constant, wherein the variation of pedestrian force with respect to spatial metrics is characterized by a power-law function, as detailed in Eq.2.
@@
$\left\{\begin{aligned} \displaystyle\Delta v&=k\Delta d^{2},\\ \displaystyle\Delta f&=m\frac{\Delta v}{\Delta t}=mk\frac{\Delta d^{2}}{\Delta t}.\end{aligned}\right$. (2)
Furthermore, in one-dimensional motion, a linear empirical relationship between the headway and speed is typically employed, often associated with the principle of constant time headway. Considering the anisotropy of the pedestrian movement space, this linear relationship, based on the nearest neighbor distance, can similarly be extended to two-dimensional motion, as illustrated in Eq.3. In this context, considering a constant relaxation time, a linear function can be employed to describe the variation of pedestrian force with respect to the relative distance to the nearest neighbor, as shown in Eq.4.
@@
$\left\{\begin{aligned} \frac{\Delta d}{\Delta v}&=t_{h},\\ \Delta f&=\frac{m\Delta v}{\Delta t}=\frac{m\Delta d}{t_{h}\,\Delta t}.\end{aligned}\right$. (4)
@@
2.2 Implicit Modeling of Anticipation and Reaction Mechanisms
In force-based methods, explicitly modeling the mechanisms of anticipation and reaction is often cumbersome. First, the underlying dynamics of the force-based model already represent the reaction mechanism through velocity adaptation constrained by a constant relaxation time. On this basis, the anticipation mechanism related to collision avoidance requires taking the future state of neighboring pedestrians (at a anticipation time of $\delta$) as input. This process potentially involves strategic interactions that require specific treatments and are therefore difficult to generalize into uniform rules. Therefore, we adopt an implicit modeling approach for anticipation mechanisms, aiming for simplicity.
@@ 2.2 Implicit Modeling of Anticipation and Reaction Mechanisms
For the simplicity, the tanh function can be employed to describe its continuous form, which aligns with the formulation used to represent the desired velocity in traffic flow. The continuous form is provided in Eq.9. Fig.4 illustrated the influence of the nearest-neighbor relative distance $d$ and angle $\theta$ on the force term.
@@
• Symmetric and asymmetric interactions
Figure 5: Schematic of nearest neighbor interaction: (a) illustration of force-based pedestrian simulation, (b) implicit modeling of anticipation and reaction mechanisms through fine-tuning in force-based models.
@@ 2.3 Model Rules
Based on the framework of a force-based method, the model is characterized by a self-driven force $\mathbf{f}_{i}$, a repulsive force $\mathbf{f}_{ij}$ exerted by the nearest entity (pedestrian or wall $j$, $j\in\mathbf{\Omega}:=\lVert{{\mathbf{d}_{ij}}}\rVert<h\wedge\angle\left({{\mathbf{v}_{i}},{\mathbf{d}_{ij}}}\right)<\phi$), and the contact forces $\mathbf{f}_{c}$ that arise during collisions. The global constraints can be expressed as follows:
$\mathbf{f}=\mathbf{f}_{i}+\mathbf{f}_{ij}+\sum_{\lVert\mathbf{d}_{ij}\rVert<r_{ij}}\mathbf{f}_{c}=m\mathbf{a}$ (10)
@@
Here, $\theta$ is the angle between the relative distance and the relative velocity, $\theta=\angle\left(\mathbf{v}_{ij},\mathbf{d}_{ij}\right)\in\left[0,\pi\right]$. Other symbols are defined as follows: $r_{ij}=r_{i}+r_{j}$, $\mathbf{v}_{ij}=\mathbf{v}_{i}-\mathbf{v}_{j}$, and $\mathbf{d}_{ij}=\mathbf{x}_{j}-\mathbf{x}_{i}$, ${\mathbf{n}_{ij}}=-{\mathbf{d}_{ij}}/\lVert{{\mathbf{d}_{ij}}}\rVert$. In this process, obstacles are modeled as stationary pedestrians and are therefore included in the calculation of the repulsive force without requiring any additions. Different from the point-to-point interaction between pedestrians, the interaction between pedestrians and obstacles follows a point-to-line relationship. Considering all potential collision scenarios, the attentional eccentricity angle $\phi$ for pedestrian-wall interaction is set to a constant value of $\pi/2$.
@@
As mentioned earlier, the pedestrian is represented as a compressible particle, described by an exponential function. When the pedestrian’s body undergoes compression, the resulting repulsive force increases exponentially with the compression magnitude. A fully repulsive core exists as the compression limit for the pedestrian particle. The parameter settings are presented in Tab.1, where $m=60$ kg, $r=0.2$ m, $\tau=0.5$ s, $t_{h}=1.3$ s, $\lambda=0.8$ and $v_{\rm{max}}$= 1.4 m/s are global parameters that remain constant across all simulations. The variable parameters $\phi$, and $\alpha$ will be adjusted as required. The frame rate for all simulations is set as 30, corresponding to a difference time of $1/30$ s.
@@
Parameter Description Type Reference value
$m$ Mass Constant 60 kg
$v_{\rm{max}}$ Maximum speed Constant 1.4 m/s
$r$ Radius Constant 0.2 m
$\tau$ Relaxation time Constant 0.5 s
$\lambda$ Dimensionless coefficient Constant 0.8 (scale in centimeters)
$t_{h}$ Approximate time headway Variable 1.3 s
$\phi$ Attentional eccentricity angle Variable [0 - $\pi$] rad
$\alpha$ Dimensionless coefficient Variable [0-1]